OpenStudy (anonymous):

limit as x approaches 0 of (sin 5x)/7x?

OpenStudy (anonymous):

I've considered multiply the numerator by 5x/5x so that sin x will become 1 as x approaches zero...but doesn't the problem then become 5x/7x and they both go to zero?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(5x)}{5x} =1\] right? so how do we get this in our problem? \[\frac{5}{7}\lim_{x \rightarrow 0}\frac{\sin(5x)}{5x}\]

OpenStudy (cathyangs):

wouldn't there be no limit? *confuzzled* O.o

myininaya (myininaya):

\[\frac{5}{7}(1)=\frac{5}{7}\]

OpenStudy (anonymous):

Sorry, is there an intermediate step missing?

OpenStudy (turingtest):

You can verify that myininaya is right very quickly if you know L'hospital's rule.

myininaya (myininaya):

all i did was multiply 5/5 @ josh